25. Then 27, n 50, 2 Suppose 20, and u = 1.96 O tH -1.645 = 3.1623 O tHo 0.0632
Answer #1

r - μΗ, tH, = -------------- (i)

Now, L = 1 , n = 50 , μΗ, = 25 and \small s^2 = 20

Substituting in (i) we get,

27 - 25 th. =-\small t_{H_{o}} = \frac{27 - 25}{\frac{\sqrt{20}}{\sqrt{50}}} = \frac{2 \times \sqrt{50}}{\sqrt{20}} = \frac{\sqrt{100}}{\sqrt{10}} = \frac{10}{\sqrt{10}} = \sqrt{10}

Therefore, th, = 3.1623